OpenStudy (anonymous):

Which number is one-fourth of the way between 4 and 12.4?

OpenStudy (anonymous):

6.1 16.4 8.4 10.3

OpenStudy (anonymous):

6.1 :)

OpenStudy (anonymous):

That's what I thought! Could you explain?

OpenStudy (zzr0ck3r):

\(4+\dfrac{1}{4}|12.4-4|\)

OpenStudy (anonymous):

8.4?

OpenStudy (zzr0ck3r):

4+0.24|8.4|=4+2.1

OpenStudy (anonymous):

6.1?

OpenStudy (anonymous):

lol sorry i was away for a second looks like @zzr0ck3r is doing the explaining for me :P

OpenStudy (anonymous):

So it is 6.1?

OpenStudy (anonymous):

I believe so yes. @zzr0ck3r Please correct me if I'm wrong.

OpenStudy (zzr0ck3r):

we both said 6.1 already...I hope that you understand why it is 6.1 then you will not need to ask if you are correct :)

OpenStudy (anonymous):

Yes i do understand i just had a spark of doubt on this one :)